To fill a water tank tap A takes 3 hours and to fill half of the tank tap B takes 2 hours. if both A and B taps are open how much time it will take to fill the tank?
Answer by stanbon(75887)

To fill a water tank tap A takes 3 hours and to fill half of the tank tap B takes 2 hours. if both A and B taps are open how much time it will take to fill the tank?
------
A DATA:
time = 3 hrs/tank ; rate = 1/3 tank/hr
---------------
B DATA:
time = 2 hrs/(1/2)tank = 4 hrs/tank ; rate = 1/4 tank/hr
-----------------------
Together DATA:
time = x hrs/tank ; rate = 1/x tank/hr
---------------------
Equation:
rate + rate = together rate
1/3 + 1/4 = 1/x
4x + 3x = 12
7x = 12
x = 12/7 = 1.714 hrs = 1hr 43 minutes (time together to fill tank )
-----
Cheers,
Stan H.
------------